# You Only Acquire Sparse-channel (YOAS): A Unified Framework for Dense-channel EEG Generation

### v1

@@ Abstract
High-precision acquisition of dense-channel electroencephalogram (EEG) signals is often impeded by the costliness and lack of portability of equipment. In contrast, generating dense-channel EEG signals effectively from sparse channels shows promise and economic viability. However, sparse-channel EEG poses challenges such as reduced spatial resolution, information loss, signal mixing, and heightened susceptibility to noise and interference. To address these challenges, we first theoretically formulate the dense-channel EEG generation problem as by optimizing a set of cross-channel EEG signal generation problems. Then, we propose the YOAS framework for generating dense-channel data from sparse-channel EEG signals. The YOAS totally consists of four sequential stages: Data Preparation, Data Preprocessing, Biased-EEG Generation, and Synthetic EEG Generation. Data Preparation and Preprocessing carefully consider the distribution of EEG electrodes and low signal-to-noise ratio problem of EEG signals. Biased-EEG Generation includes sub-modules of BiasEEGGanFormer and BiasEEGDiffFormer, which facilitate long-term feature extraction with attention and generate signals by combining electrode position alignment with diffusion model, respectively. Synthetic EEG Generation synthesizes the final signals, employing a deduction paradigm for multi-channel EEG generation. Extensive experiments confirmed YOAS’s feasibility, efficiency, and theoretical validity, even remarkably enhancing data discernibility. This breakthrough in dense-channel EEG signal generation from sparse-channel data opens new avenues for exploration in EEG signal processing and application.
@@
Multi-channel, Sparse-channel, Dense-channel, EEG signal generation, Transformer, Generative adversarial network, Diffusion model, Neuroscience, Brain-computer interface
@@ I Introduction
Electroencephalogram (EEG), crucial for neuroscience and clinical diagnosis, captures neuronal electrical activity from the cerebral cortex. In research, EEG offers new insights into brain function, elucidating cognitive processes, sleep patterns, neural mechanisms of brain disorders, etc [1, 2, 3, 4]. In clinical settings, EEG aids in diagnosing epilepsy [5], sleep disorders [6], and informs surgical decisions [7]. It also can assesse treatment efficacy [8] and contribute to advancements in brain-computer interface (BCI) technology [9], enhancing rehabilitation and quality of life for individuals with disabilities.
@@
Neural activity is influenced by factors such as neuron types, synaptic connections, and neural circuits [10]. And as a technique for exploring neural activity, EEG are high-dimensional and dynamic, necessitating the capture of intricate temporal correlations and spatial distributions [11, 12]. Therefore, generating EEG signals presents challenges in simulating the complex brain activity, which is characterized by the spatiotemporal properties of each channel. Moreover, generated EEG signals exhibit uncertainty and variability due to physiological and environmental factors, posing limitations to model training and performance evaluation [13, 14]. In summary, the generation of multi-channel EEG signals faces the following challenges:
Firstly, as each electrode corresponds to specific brain areas, harmonizing signal characteristics from electrodes at different locations under the same mental condition presents challenges. This is due to that the distinct characteristics of EEG signals vary among individuals across different inner states [15], and the electrode signals exhibit unique features at various positions [16, 17].
Secondly, EEG signal generation for whole brain only based on few channels is exceptionally challenging. Initially, the low spatial resolution of EEG signals makes pinpointing their sources quite difficult. Further, within the International 10-20 system, the electrode signals in the neighboring areas may have weak correlation, while those from distant areas may exhibit strong correlation. Additionally, the previous studies [18] have only focused on generating EEG data tied to specific stimuli.
Thirdly, the generation models of EEG signals pose the challenges including training instability, inefficiency, and susceptibility to signal noise. Despite the advancement of generation models such as Variational Auto Encoder (VAE) [19] and Generative Adversarial Network (GAN) series [20] in general domain, these issues persist for the specific EEG signal generation problem.
Figure 1: The YOAS framework within the exploratory scheme includes Data Preparation, extracting Bias Signal from Raw Data through regional division, followed by Data Preprocessing to eliminate outliers. The Biased-EEG Generation module captures long-term EEG features and aligns spatial positions to generate Two-stage Bias. Finally, the Synthetic EEG Generation module synthesizes the final EEG signal.
Therefore, to address the aforementioned challenges, we proposed a specialized model for dense-channel EEG generation problem, named YOAS comprising four sequential modules, depicted in Fig. 1, to improve the efficiency of EEG signal generation for whole brain:
@@
Data Preprocessing: In the context of EEG bias signals derived from the Data Preparation stage, we commence by empirically eliminating data points within each signal segment deviating by one or two standard deviations from the total segment signal of each channel. Subsequently, linear interpolation is employed to address the missing data points. Finally, Multiscale Principal Components Analysis (MSPCA) [23] is utilized to reduce dimensionality and extract features from the interpolated data.
Biased-EEG Generation: The Biased-EEG Generation modules comprise BiasEEGGanFormer and BiasEEGDiffFormer. BiasEEGGanFormer captures high timing resolution characteristics using Transformer [24] with multi-head attention, producing the “One-stage Generated Bias”. In BiasEEGDiffFormer, spatial alignment and feature extraction are performed using ConvBlock. DiffusionBlock then generates signal samples incrementally from noise data, ensuring consistency with electrode locations, resulting in the “Two-stage Generated Bias”.
Synthetic EEG Generation: In this module, we synchronize the Regional Reference Signal with the Two-stage Generated Bias to achieve spatial alignment. Then, we add the Regional Reference Signal to the Two-stage Generated Bias to generate signals for each channel. To cover the entire brain, we adopt divide-and-conquer approach, selecting key channels within each region as references and optimizing the derivation path with greedy algorithm.
@@
Formulation of the Dense-channel EEG Generation Problem: We formulate the problem of generating dense-channel EEG signals from sparse channels as a cross-channel EEG generation task. This formulation addresses challenges related to biased EEG signal generation, EEG electrode position alignment and the synthesis of EEG signals across channels.
@@ II-A Across-channel EEG Signal Generation
In this study, we delineate the cross-channel EEG signal generation task as the process of generating an additional EEG channel signal based on the collected EEG channel signal. Firstly, previous studies predominantly concentrated on data augmentation, but failed to accomplish cross-channel EEG signal generation. For instance, EEG-GAN [20] primarily targets the single-channel data augmentation, showing an enhancement over WGAN [25]; Nonetheless, its pursuit of capturing features in lengthy sequences and local intricacies escalates computational costs and training instability significantly. Moreover, although BioDiffusion [26] excels in producing high-fidelity multivariate signals, it does not specifically cater to cross-channel EEG signal generation. Consequently, we introduce the Biased-EEG Generation module incorporating Transformer [24] and diffusion model [27, 28], thereby facilitating cross-channel EEG signal generation.
@@ II-B EEG Electrode Position Alignment
The EEG electrode position alignment refers to the process of aligning generated EEG signal data with the physical positions of the electrodes to which they correspond. Generally speaking, diffusion models [27] operating by incrementally introducing and reducing noise through both forward and reverse processes, is widely used to generate the data in specific conditions. The introduction of DDPM [28] further enhances the model, surpassing GAN [29] in generating high-quality images. While diffusion models have been applied to generate high-fidelity, non-stationary biomedical signals [26], they primarily focus on enhancing single-channel data, not multi-channel or dense-channel data. Thus, despite the advancements, a crucial issue that remains unresolved is the lack of adherence to the positional characteristics of the electrodes in the generated signals. In BiasEEGDiffFormer of YOAS, we propose an innovative diffusion model to align electrode positions by matching signals generated by BiasEEGGanFormer with corresponding reference signals in their respectively divided regions.
@@ II-C Across-channel EEG Synthetic Paradigm
The across-channel EEG synthetic paradigm refers to the deduction paradigm process that pursuing the optimized sets of the generation paths for generating multi-channel EEG signals across the whole brain. As it stated before, most of EEG signal generation studies primarily focused on signal generation for data augmentation [10, 26]. However, in the pursuit of generating multi-channel EEG signals throughout the entire brain, the task of deriving EEG signals for all channels from a single channel signal frequently presents significant challenges. Hence, it becomes imperative to determine and refine the optimal number of reference channel signals required to generate signals across all channels. Additionally, the optimization and delineation of generation pathways for specific channel signals based on the designated reference channel signals are also paramount. In response to these challenges, we introduce the Synthetic EEG Generation module within the YOAS framework to effectively address the aforementioned concerns.
@@ III-A Overview
The task for dense-channel EEG signal generation for entire brain using YOAS, as depicted in Fig. 1, could be theoretically formulated by optimizing a set of across-channel EEG signal generation problems based on $K$ reference channels, where each reference channel represents a regional division (RD). Initially, as shown in Data Preparation stage of Fig. 1, the entire brain positions are partitioned into $K$ RDs according to the widely-used International 10-20 system in EEG, denoted as $RD_{i}$ for the $i$-th division and $M_{i}$ as the total number of channels within $RD_{i}$. Secondly, we can record the channel number of the $i$-th RD as the set $\hat{M}_{i}$, $\hat{M}_{i}=\{1,2,\ldots,j,\ldots,M_{i}\}$, with $j$ standing for the $j$-th channel ($1<j\leq M_{i}$) in the $RD_{i}$. Thirdly, the whole brain channel signal set can be formulated as $\{O_{ij}\}_{i=1,j=1}^{K,M_{i}}$, the reference channel signal set can be defined as $\{O_{i1}\}_{i=1}^{K}$ by selecting the first channel for each RD and the remaining channel signal set can be represented by $\{O_{ij}\}_{i=1,j=2}^{K,M_{i}}$. Thus, the across-channel EEG signal generation problem is involved by asking that “can the remaining channel signals $O_{ij}$ ($1<j\leq M_{i}$) in the $RD_{i}$ be accurately generated based on the reference channel $O_{i1}$, subject to $0<i\leq K$?” Mathematically, the dense-channel EEG signal generation problem of whole brain can be initially expressed as follows:
@@ III-B Data Preparation
This section elucidates the initial regional division discussed in Section III-A, which divides all brain channels into $K$ RDs using the International 10–20 system of EEG electrode arrangement and the approximately structural/functional symmetry between the left and right hemispheres [16]. To achieve this, we propose three hypotheses for multi-channel EEG generation pathways, illustrated in Fig. 2. For convenience, we denote that $O_{ij}$ represents the source channel, and $O_{im}$ represents the target channel within the $RD_{i}$. Moreover, $Dis(.)$ calculates the physical distance between two electrodes, while $D(.)$ computes the correlation distance between their signals, where $D=\mid 1-\rho\mid$ $(-1<\rho<1)$, with $\rho$ representing the Pearson Correlation Coefficient.
Hypothesis 1. In Fig. 2(A), the signal of a given channel may be generated by another nearby channel. If the criteria
@@
$\begin{cases}Dis(O_{ij},O_{im})\leq L_{1}\\ D(O_{ij},C_{im})\leq P_{1}\end{cases}$ is met, it implies that the signal of the target channel $O_{im}$ can be generated through the source channel $O_{ij}$, i.e., $O_{ij}\longrightarrow O_{im}$.
Hypothesis 2. In Fig. 2(B), for a specific channel, its signal may be indirectly derived from another channel within the same $RD_{i}$ through intermediate channel $O_{in}$. If the condition
@@
$\begin{cases}\begin{cases}Dis(O_{ij},O_{in})\leq L_{2}\&Dis(O_{ij},O_{im})\leq L_{2}\\ \&Dis(O_{im},O_{in})\leq L_{2}\end{cases}\\ D(O_{ij},C_{in})\leq P_{2}\\ D(O_{in},C_{im})\leq P_{2}\end{cases}$ is satisfied, it indicates that the signal of target channel $O_{im}$ can be indirectly generated by that of the source channel ${O_{ij}}$, i.e., $O_{ij}\xrightarrow{o_{in}}O_{im}$.
@@
Hypothesis 3. In Fig. 2(C), for channels with certain level of symmetry between the left and right hemispheres of the brain, their signals can mutually generate. Thus, if the criteria $\begin{cases}Dis(O_{ij},O_{im})\leq L_{3}\\ D(O_{ij},C_{im})\leq P_{3}\\ D(C_{ij},O_{im})\leq P_{3}\end{cases}$ is satisfied, this signifies a mutual signal generation between the source channel $O_{ij}$ and the target channel $O_{im}$, i.e., $O_{ij}\longleftrightarrow O_{im}$.
@@
To this end, we define the maximum physical radiation generation range set of the reference channel in the above three hypotheses as $L=\{L_{1},L_{2},L_{3}\}$, and the coefficient set of the maximum feasible generation signal of the reference channel signal is recorded as $P=\{P_{1},P_{2},P_{3}\}$, $\left.\left(\begin{matrix}0\leq P_{i}\leq 2\\ \end{matrix}\right.\right)$. In the International 10-20 system for EEG placement, zoning proceeds as follows: initially, we empirically partition it into 5 regions comprising the Frontal Lobe, Temporal Lobe, Central Sulcus, Parietal Lobe, and Occipital Lobe. Subsequently, leveraging the approximately hemispherical symmetry, we utilize the central sulcus as the axis for division, excluding paired regions with notable signal correlation such as (A1, A2) and (O1, O2). Further, we apply the aforementioned hypotheses to refine the regional division task. Hypothesis 1 (III-B) sets $L_{1}$ empirically to the radius of the 2D distribution map with $P_{1}$ equal to 0.3. Hypothesis 2 (III-B) maintains $L_{2}$ at the 2D distribution map’s radius, with $P_{2}$ equal to 0.3. Hypothesis 3 (III-B) sets $L_{3}$ to the 2D distribution map’s diameter, with $P_{3}$ equal to 0.1.

### v2

@@ Abstract
High-precision acquisition of dense-channel electroencephalogram (EEG) signals is often impeded by the costliness and lack of good portability. In contrast, generating dense-channel EEG signals from sparse channels shows promise and economic viability. However, sparse-channel EEG poses challenges such as reduced spatial resolution, information loss, signal mixing, and heightened susceptibility to noise and interference. To address these challenges, we first theoretically formulate the dense-channel EEG generation problem as by optimizing a set of cross-channel EEG signal generation problems. Then, we propose the YOAS framework for generating dense-channel data from sparse-channel EEG signals. The YOAS totally consists of four sequential stages: Data Preparation, Data Preprocessing, Biased-EEG Generation, and Synthetic EEG Generation. Data Preparation and Preprocessing carefully consider the distribution of EEG electrodes and low signal-to-noise ratio problem of EEG signals. Biased-EEG Generation includes sub-modules of BiasEEGGanFormer and BiasEEGDiffFormer, which facilitate long-term feature extraction with attention and generate signals by combining electrode position alignment with diffusion model, respectively. Synthetic EEG Generation synthesizes the final signals, employing a deduction paradigm for multi-channel EEG generation. Extensive experiments confirmed YOAS’s feasibility, efficiency, and theoretical validity, even remarkably enhancing data discernibility. This breakthrough in dense-channel EEG signal generation from sparse-channel data opens new avenues for exploration in EEG signal processing and application.
@@
Sparse-channel, Dense-channel, EEG signal generation, GAN, Diffusion model, BCI
@@ I Introduction
Electroencephalogram (EEG), crucial for neuroscience and clinical diagnosis, captures neuronal electrical activity from the cerebral cortex[1]. In academical research, EEG offers new insights into brain function, elucidating cognitive processes, sleep patterns, neural mechanisms of brain disorders, etc [2], [3], [4], [5], [6], [7]. In clinical settings, EEG aids in diagnosing epilepsy [8], [9], sleep disorders [10], [11], [12], and informs surgical decisions [13], [14]. It also can assesse treatment efficacy [15] and contribute to advancements in brain-computer interface (BCI) technology [16], [17], [18], [19], [20], enhancing rehabilitation and quality of life for individuals with disabilities.
@@
Neural activity is influenced by factors such as neuron types, synaptic connections, and neural circuits [21]. EEG signals recording neural activity are high-dimensional and dynamic, necessitating the capture of intricate temporal correlations and spatial distributions [22], [23]. Thus, generating EEG signals presents challenges in simulating the spatiotemporal properties of each channel. Moreover, generated EEG signals exhibit uncertainty and variability due to physiological and environmental factors, posing limitations to model training and performance evaluation [24], [25]. In summary, the generation of dense-channel EEG signals faces the following challenges:
Firstly, as each electrode corresponds to specific brain areas, harmonizing signal characteristics from electrodes at different locations under the same mental condition presents challenges. This is due to that the distinct characteristics of EEG signals vary among individuals across different inner states [26], [27], and the electrode signals exhibit unique features at various positions [28], [29].
Secondly, EEG signal generation for whole brain based on few channels is exceptionally challenging. Initially, the low spatial resolution of EEG makes pinpointing their sources quite difficult. Further, within the International 10-20 system, the channel signals in the neighboring areas may have weak correlation, while those from distant areas may exhibit strong correlation. Additionally, the previous studies [30] have only focused on generating EEG data as augmentation.
Thirdly, the generative models of EEG signals pose the challenges including training instability, inefficiency, and susceptibility to signal noise. Despite the advancement of generative models such as Variational Auto Encoder (VAE) [31], Generative Adversarial Network (GAN) series [32], and Diffusion models [33], [34] in general domain, these issues persist for the specific EEG signal generation problem.
Figure 1: The YOAS framework. The Data Preparation module extracts Bias Signal from Raw Data through regional division, followed by Data Preprocessing to eliminate outliers. The Biased-EEG Generation module captures long-term features and aligns spatial positions to generate Two-stage Bias. Finally, the Synthetic EEG Generation module synthesizes the final signals.
To address the aforementioned challenges, we proposed a specialized model for dense-channel EEG generation problem, named YOAS comprising four sequential modules, depicted in Fig. 1, to improve the efficiency of EEG signal generation:
@@
Data Preprocessing: Based on the EEG bias signals derived from Data Preparation module, we commence by empirically eliminating data points within each signal segment deviating by one or two standard deviations from the total segment signal of each channel. Subsequently, linear interpolation is employed to address the missing data points. Finally, Multiscale Principal Components Analysis (MSPCA) [37] is utilized to reduce dimensionality and extract features from the interpolated data.
Biased-EEG Generation: The Biased-EEG Generation module comprises BiasEEGGanFormer and BiasEEGDiffFormer. BiasEEGGanFormer captures high timing resolution characteristics using Transformer [38] with multi-head attention, producing the “One-stage Generated Bias”. In BiasEEGDiffFormer, spatial alignment and feature extraction are first performed by ConvBlock, followed by DiffusionBlock that generates signal samples incrementally from noise data, ensuring consistency with electrode locations, resulting in the “Two-stage Generated Bias”.
Synthetic EEG Generation: In this module, we add the Regional Reference Signal to the Two-stage Generated Bias to generate signals for each channel. Specifically, we adopt divide-and-conquer approach, selecting key channels within each region as references and optimizing the derivation path with greedy algorithm.
@@
Formulation of the Dense-channel EEG Generation Problem: We formulate the problem of generating dense-channel EEG signals from sparse channels for whole brain as a set of cross-channel EEG generation tasks. This formulation addresses challenges related to biased EEG signal generation, EEG electrode position alignment and the synthesis of EEG signals across channels.
@@ II-A Across-channel EEG Signal Generation
In this study, we delineate the cross-channel EEG signal generation task as the process of generating an additional EEG channel signal based on the collected EEG channel signals. Firstly, previous studies predominantly concentrated on data augmentation, but failed to accomplish cross-channel EEG signal generation. For instance, EEG-GAN [32] primarily targets the single-channel data augmentation, showing an enhancement over WGAN [39]; Nonetheless, its pursuit of capturing features in lengthy sequences and local intricacies escalates computational costs and training instability significantly. Moreover, although BioDiffusion [40] excels in producing high-fidelity multivariate signals, it does not specifically cater to cross-channel EEG signal generation. Consequently, we introduce the Biased-EEG Generation module incorporating Transformer [38] and diffusion model [41], [42], thereby facilitating cross-channel EEG signal generation.
@@ II-B EEG Electrode Position Alignment
The EEG electrode position alignment refers to the process of aligning generated EEG signal data with the physical positions of the electrodes to which they correspond. Generally speaking, diffusion models [41] operating by incrementally introducing and reducing noise through both forward and reverse processes, is widely used to generate the data. The proposed DDPM [42] further enhances the model, surpassing GAN [43] in generating high-quality images. While diffusion models have been applied to generate high-fidelity, non-stationary biomedical signals [40], they primarily focus on enhancing single-channel data, not multi-channel or cross-channel data. Thus, a crucial issue that remains unresolved is the lack of adherence to the positional characteristics of the electrodes in the generated signals. In BiasEEGDiffFormer of YOAS, we propose an innovative diffusion model to align electrode positions by matching signals generated by BiasEEGGanFormer with corresponding reference signals in their respectively divided regions.
@@ II-C Across-channel EEG Synthetic Paradigm
The across-channel EEG synthetic paradigm refers to the deduction paradigm process that pursuing the optimized sets of the generation paths when generating multi-channel EEG signals across the whole brain. Most of EEG signal generation studies primarily focused on signal generation for data augmentation [21], [40]. However, in the pursuit of generating multi-channel EEG signals throughout the entire brain, the task of deriving EEG signals for all channels from a single channel signal frequently presents significant challenges. Hence, it becomes imperative to determine and refine the optimal number of reference channel signals required to generate signals across all channels. Additionally, the optimization and delineation of generation pathways for specific channels based on the designated reference channels are also paramount. In response to these challenges, we introduce the Synthetic EEG Generation module within the YOAS framework to effectively address the aforementioned concerns.
@@ III-A Overview
The task for dense-channel EEG signal generation for entire brain using YOAS, as depicted in Fig. 1, could be theoretically formulated by optimizing a set of across-channel EEG signal generation problems based on $K$ reference channels, where each reference channel represents a regional division (RD). Initially, as shown in Data Preparation of Fig. 1, the entire brain positions are partitioned into $K$ RDs according to the widely-used International 10-20 system in EEG, denoted as $RD_{i}$ for the $i$-th division and $M_{i}$ as the total number of channels within $RD_{i}$. Secondly, we can record the channel index within $RD_{i}$ as set $\hat{M}_{i}=\{1,2,\ldots,j,\ldots,M_{i}\}$, with $j$ standing for the $j$-th channel ($1<j\leq M_{i}$) of $RD_{i}$. Thirdly, the whole brain channel signal set can be formulated as $\{O_{ij}\}_{i=1,j=1}^{K,M_{i}}$, the reference channel signal set can be defined as $\{O_{i1}\}_{i=1}^{K}$ by selecting the first channel for each RD, and the remaining channel signal set can be represented by $\{O_{ij}\}_{i=1,j=2}^{K,M_{i}}$. Thus, the across-channel EEG signal generation problem is involved by asking that “Can the remaining channel signals $O_{ij}$ ($1<j\leq M_{i}$) in the $RD_{i}$ be accurately generated based on the reference channel $O_{i1}$, subject to $0<i\leq K$?” Mathematically, the dense-channel EEG signal generation problem of whole brain can be initially expressed as follows:
@@ III-B Data Preparation
This section elucidates the initial regional division discussed in Section III-A, which divides all brain channels into $K$ RDs based on the International 10–20 system and the approximately structural/functional symmetry between the left and right hemispheres [28]. To achieve this, we propose three hypotheses for multi-channel EEG generation pathways, illustrated in Fig. 2. For convenience, we denote that $O_{ij}$ represents the source channel, and $O_{im}$ represents the target channel within the $RD_{i}$. Moreover, $Dis(.)$ calculates the physical distance between two electrodes, while $D(.)$ computes the correlation distance between their signals, where $D=\mid 1-\rho\mid$ $(-1<\rho<1)$, with $\rho$ representing the Pearson Correlation Coefficient.
Hypothesis 1. In Fig. 2(A), if the criteria
@@
$\begin{cases}Dis(O_{ij},O_{im})\leq L_{1}\\ D(O_{ij},C_{im})\leq P_{1}\end{cases}$ is met, it implies that the signal of the target channel $O_{im}$ can be directly generated through the source channel $O_{ij}$, i.e., $O_{ij}\longrightarrow O_{im}$.
Hypothesis 2. In Fig. 2(B), for a specific channel, if the condition
@@
$\begin{cases}\begin{cases}Dis(O_{ij},O_{in})\leq L_{2}\&Dis(O_{ij},O_{im})\leq L_{2}\\ \&Dis(O_{im},O_{in})\leq L_{2}\end{cases}\\ D(O_{ij},C_{in})\leq P_{2}\\ D(O_{in},C_{im})\leq P_{2}\end{cases}$ is satisfied, it indicates that the signal of target channel $O_{im}$ can be indirectly generated by that of the source channel ${O_{ij}}$, i.e., $O_{ij}\xrightarrow{o_{in}}O_{im}$ through intermediate channel $O_{in}$.
@@
Hypothesis 3. In Fig. 2(C), for channels demonstrating strong symmetry between the left and right hemispheres, if the criteria $\begin{cases}Dis(O_{ij},O_{im})\leq L_{3}\\ D(O_{ij},C_{im})\leq P_{3}\\ D(C_{ij},O_{im})\leq P_{3}\end{cases}$ is satisfied, this signifies a mutual signal generation between the source channel $O_{ij}$ and the target channel $O_{im}$, i.e., $O_{ij}\longleftrightarrow O_{im}$.
@@
To this end, we record the maximum physical radiation generation range set and the coefficient set in the aforementioned three hypotheses as $L=\{L_{1},L_{2},L_{3}\}$ and $P=\{P_{1},P_{2},P_{3}\}$ $\left.\left(\begin{matrix}0\leq P_{i}\leq 2\\ \end{matrix}\right.\right)$, respectively. In the International 10-20 system for EEG placement, zoning procedure proceeds as follows: initially, we empirically partition it into 5 regions comprising the Frontal Lobe, Temporal Lobe, Central Sulcus, Parietal Lobe, and Occipital Lobe. Subsequently, leveraging the approximately hemispherical symmetry, we utilize the central sulcus as the axis for division, excluding paired regions with notable signal correlation such as (A1, A2) and (O1, O2). Further, we apply the aforementioned hypotheses to refine the regional division task. Hypothesis 1 (III-B) sets $L_{1}$ empirically to the radius of the 2D distribution map with $P_{1}$ equal to 0.3. Hypothesis 2 (III-B) maintains $L_{2}$ at the 2D distribution map’s radius, with $P_{2}$ equal to 0.3. Hypothesis 3 (III-B) sets $L_{3}$ to the 2D distribution map’s diameter, with $P_{3}$ equal to 0.1.